Find the number of distinguishable permutations of the given letters "AABCD
Answer #1

Total number of permutations = 5!

Repeatedly occurinoc word adjusted = 2!

Hence, total number of distinguishable permutations = 5!/2!

= 1*2*3*4*5/1*2 = 3*4*5=60

Hence, total number of distinguishable permutations = 60